In a hypothetical economy, without taxes, the consumption(C) at different levels of real GDP (Y) is
given here under:
Real GDP Consumption (Rs in Billion)
(Y) (C)
3,000 1,500
5,000 3,000
Determine the size of MPC (Marginal propensity to consume), MPS (Marginal propensity to save) and multiplier
1
Expert's answer
2020-09-16T20:58:50-0400

"MPC = \dfrac{ \Delta C} { \Delta Y}
 = \dfrac {3000-1500} {5000-3000}
 = 0.75"

"MPS = 1 - MPC\n = 1 - 0.75\n = 0.25"

"Multiplier = {1 \above{1pt} 1- MPC}
 = {1 \above{1pt} 1- 0.75}
 = 4"